OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the following circle. (x + 2)^2 + y^2 = 10 center= radius= sqrt ?

OpenStudy (hoblos):

(-2,0) sqrt10

OpenStudy (anonymous):

(-2,0) \[r= \sqrt{10}\]

OpenStudy (ash2326):

compare with the standard equation to find the center and radius

OpenStudy (anonymous):

standard equation is (x-a)^2 + (y-b)^2 = r^2 (a,b) is center r = radius

OpenStudy (anonymous):

Are you getting the hang of this type of question?